# PIPNet3D: Interpretable Detection of Alzheimer in MRI Scans

### v1

@@ 4 Conclusion
In summary, we introduced PIPNet3D, a part-prototype model for 3D input images, applied to diagnose AD from sMRI, whose interpretability has been systematically assessed. We obtained an interpretable and compact model able to support Alzheimer’s diagnosis with prototypes in line with medical knowledge. Our results show, that interpretable models that align with domain-knowledge can be as effective as blackbox models. We hope to inspire their usage in the future, especially in high-risk domains, such as medical image analysis. We plan to extend PIPNet3D to a multi-modal system additionally taking patients’ demographic information as input, and to increase the number of predicted classes by introducing intermediate levels of cognitive impairments. We also plan to design experiments to evaluate OoD performances with intermediate levels of dementia and patients with different age groups.

### v3

@@ 4 Conclusion
In summary, we introduced PIPNet3D, a part-prototype model for 3D input images, applied to diagnose AD from sMRI, whose interpretability has been systematically assessed. We obtained an interpretable and compact model able to support AD diagnosis with prototypes in line with medical knowledge. Our results show, that interpretable models that align with domain-knowledge can be as effective as blackbox models. We hope to inspire their usage in the future, especially in high-risk domains, such as medical image analysis. We plan to extend PIPNet3D to a multi-modal system additionally taking patients’ demographic information, introducing intermediate levels of cognitive impairments, and evaluating OoD performances with intermediate levels of dementia and patients with different age groups.